# Breaking the Script Barrier in Multilingual Pre-Trained Language Models with Transliteration-Based Post-Training Alignment

### v1

@@ 1 Introduction
Many post-training alignment strategies use objectives that rely on bilingual dictionaries or parallel data to align the representations of mPLMs (Cao et al., 2020; Wang et al., 2020; Schuster et al., 2019; Pan et al., 2021). However, dictionaries and parallel corpora are often limited in the data scale or the number of languages they cover (Artetxe et al., 2020b), which might be impractical for building strong supervision signals for many low-resource languages. Another alternative that improves cross-lingual alignment is to use transliteration (a process of converting the text of a language from one script to another). Transliteration can improve the lexical overlap, especially for related languages (Moosa et al., 2023). Different from translation, transliterations can be obtained nearly for free using well-performing rule-based tools (Hermjakob et al., 2018). Therefore, several works have shown improvements in cross-lingual transfer by pre-training or fine-tuning models with data transliterated into a common script (Murikinati et al., 2020; Muller et al., 2021; Purkayastha et al., 2023; Moosa et al., 2023). However, they require the use of a single common script by the model. This is restrictive for many tasks as the process of transliteration can be lossy and non-invertible.
@@
Recently, Liu et al. (2024a) proposed a sequence-level contrastive learning objective to improve the alignment across different scripts at a large scale (for more than 500 languages), using sentences in both their original scripts and their Latin-script transliterations, validated by English-centric cross-lingual zero-shot transfer evaluations. However, there are three major limitations in their setup. First, the contrastive objective only manipulates the sequence-level representations in the middle layer, which does not directly contribute to better alignment in the token-level space. Second, not every language pair has extensive lexical overlap that can boost cross-lingual transfer: transliteration-based alignment makes more sense for mutually influenced languages. Lastly, English alone as a transfer source language does not fully exploit the alignment benefit, as it does not have the most lexical overlap with other languages.
@@
To this end, we propose a new transliteration-based post-training alignment method that works on both sequence and token levels. Our method does not rely on parallel data. Instead, similar to Liu et al. (2024a), we use the monolingual data in their original scripts and their Latin transliteration obtained by using Uroman (Hermjakob et al., 2018), a rule-based transliteration tool. We investigate the impact of the strategy by focusing on two groups of languages: Mediterranean-Amharic-Farsi and South+East Asian Languages, described more in detail in Section 4.2. The languages in each group share areal features but differ in scripts. Some languages are closely related as members of the same language family (e.g., Semitic and Sino-Tibetan). Additionally, languages in each group have extensive lexical overlap due to historical contact and geographical proximity (e.g., Chinese and Korean or Turkish and Arabic). As these languages are written in different scripts, transliteration can help to better exploit the shared linguistic properties and thus improve the cross-lingual transfer.
@@
We leverage our method to post-train Glot500 (ImaniGooghari et al., 2023) (a continually pre-trained model from XLM-R on more than 500 languages) on the selected language groups and evaluate the zero-shot cross-lingual transfer performance on three types of downstream tasks: sentence retrieval, text classification, and sequence labeling. The evaluation is done with English as the source language, as well as with three other source languages of different scripts in each group. We show that our method consistently improves the downstream task performance across different languages and scripts. Moreover, our method further boosts the transfer performance when better source languages are chosen, as the performance depends on the degree of alignment between the source and target languages – precisely the alignment that our approach boosts.
@@ 2 Related Work
Many recent works have proposed pre-training or fine-tuning alignment methods to improve cross-lingual transfer in mPLMs. Cao et al. (2020) proposed a fine-tuning embedding alignment objective between word pairs procured in an unsupervised fashion from parallel data using statistical word alignment models (Dyer et al., 2013). Chaudhary et al. (2020) improved alignment during pre-training by using bilingual dictionaries to replace words in original sentences with translations in other languages. Similarly, Tang et al. (2022) used bilingual dictionaries to explicitly align the embeddings of the same words in different languages during pre-training. Wei et al. (2021) proposed a hierarchical contrastive learning pre-training method, which uses parallel data to align representations at the word and sentence levels. Similarly, Hu et al. (2021) proposed a pre-training method with explicit alignment signals from parallel data that encourages symmetry at both word and sentence levels. Pan et al. (2021) combined contrastive learning with translation language modeling (Conneau and Lample, 2019) as a post-training alignment method that uses parallel data as well. While these methods have shown improvements in cross-lingual transfer, they have the limitation of requiring parallel data or bilingual dictionaries, which may be hard to acquire for many low-resource languages.
@@
Transliteration is a process of converting the text of a language from one script to another (Wellisch et al., 1978). This process does not involve translating meanings but rather represents the original symbols as closely as possible in the target script. Different works have proposed transliteration-based methods to address the script barrier problem in multilingual models. Murikinati et al. (2020) used transliteration to a common script to improve cross-lingual morphological inflection. Khemchandani et al. (2021) exploited language relatedness between Indian languages and leveraged transliteration to a common script to adapt multilingual models to low-resource languages. Muller et al. (2021) analyzed the behavior of multilingual models on unseen languages and found that languages written in different scripts do not benefit from transfer learning. They proposed transliteration to the high-resource source language script as a solution to address the script barrier. Purkayastha et al. (2023) showed that fine-tuning multilingual models on data transliterated into Latin script improves cross-lingual transfer for low-resource languages. Similarly, Moosa et al. (2023) pre-trained models from scratch on data transliterated into a common script for the Indic languages and showed improvements in cross-lingual transfer. Our work is most related to TransliCo proposed by Liu et al. (2024a), where a sequence-level contrastive learning objective is used to encourage alignment across different scripts without restricting the models to a common script, using data in their original scripts and their Latin-script transliteration. However, their English-centric evaluation setup limits the ability to fully reveal the impact of transliteration. This paper systematically explores how transliteration-based alignment enhances the performance of using various source languages.
@@ 3 Methods
We present a transliteration-based post-training alignment method that can be used to fine-tune existing encoder-only mPLMs for improved alignment across languages using different scripts, boosting cross-lingual transfer performance. Our method consists of three objectives: masked language modeling, sentence-level alignment, and token-level alignment. All objectives are trained on combined original and transliterated data. The transliterated data is obtained by converting the original data into Latin script. The transliteration process uses Uroman (Hermjakob et al., 2018), a rule-based system that can convert nearly all character sets into a common Latin script. The overall method is illustrated in Figure 1, and we introduce the three objectives in detail in the following.
@@ 3.1 Masked Language Modeling
where $\mathcal{M}$ is the set of masked positions in the input sentence $X_{i}$ (either $X^{orig}_{i}$ or $X^{latn}_{i}$) and $P_{\text{MLM}}(X_{i,m}|\boldsymbol{h}_{i,m})$ is the probability of predicting token $X_{i,m}$ giving $\boldsymbol{h}_{i,m}$, the final contextualized representation at the position $m$ in the $i$th sequence. The probability is computed by an MLM head. Fine-tuning with MLM on original data is necessary to preserve the model’s knowledge. On the other hand, the mPLM has very limited knowledge about the transliterated data, which makes the MLM objective on transliterated data crucial for learning useful cross-script representations. We refer to the MLM objective for the original data (resp. transliterated data) as $\mathcal{L}_{\text{MLM}}^{orig}$ (resp. $\mathcal{L}_{\text{MLM}}^{latn}$).
@@ 3.2 Sentence-Level Alignment
where $\text{NEG}=\sum_{(X,X^{-})\in B^{-}}e^{\text{sim}(f(X),f(X^{-}))/\tau}$, $f$ is defined as mean pooling over the 8th layer output contextualized embeddings (ignoring the special tokens output except for [mask] token), sim is the dot product, and $\tau$ is the temperature set to $1$.
@@ 3.3 Token-Level Alignment
The sentence-level alignment objective helps the model to learn similar sentence representations for the original and transliterated sequences. This is useful for improving performance in sentence-level downstream tasks like sentence retrieval or classification. However, this objective manipulates the output of a middle layer, which does not directly contribute to better alignment in the token-level space. For token-level tasks like NER and POS tagging, alignment at the token level might be more beneficial. Therefore, we propose a token-level alignment objective that further encourages the model to align the representations of the original and transliterated words. We adapt the translation language modeling objective introduced by  Conneau and Lample (2019), which is equivalent to applying the MLM objective on a concatenated bilingual sentence pair. Specifically, given a sentence pair $(X^{orig}_{i},X^{latn}_{i})$, we apply the MLM objective on the concatenated sequence $X^{orig}_{i}\oplus X^{latn}_{i}$ or $X^{latn}_{i}\oplus X^{orig}_{i}$, where concatenation order is randomly chosen during training. The intuition is that, to predict a token masked in the original sentence, the model can either attend to surrounding tokens in the original script or their transliterations and vice versa. This encourages the model to align the representations in the original script and the Latin script. We refer to this objective as transliteration language modeling (TLM) and the loss as $\mathcal{L}_{\text{TLM}}$.
@@ 4.2 Languages, Data and Models
The two language groups are named as Mediterranean-Amharic-Farsi and South+East Asian Languages. We visualize each group’s geographical distribution of the selected languages in Figure 2. The languages within each group are spoken in adjacent areas and there is a long history of linguistic influence between them. For example, Arabic has had extensive contact with languages such as Turkish and Persian (Versteegh, 2001). The data for each language is sampled from the Glot500-c training dataset (ImaniGooghari et al., 2023). We sample 10% of the available data for each language, or a minimum of 10k sentences, whichever is larger. The data is then transliterated into Latin script using Uroman (Hermjakob et al., 2018). Table 1 shows each group’s languages and the number of sampled sentences. In total, Mediterranean-Amharic-Farsi consists of 10 languages, 5 scripts and around 16M sentences, while South+East Asian Languages consists of 10 languages, 7 scripts and around 4M sentences. We fine-tune Glot500 using our alignment method on each group separately. We then select the best checkpoint for each group by validating the checkpoints’ performance on the Tatoeba (Artetxe and Schwenk, 2019) sentence retrieval dataset, which contains 1000 English-aligned sentences. We compute the top-10 retrieval accuracy based on the cosine similarity of the averaged 8th-layer contextual embeddings. The best checkpoint for each group is regarded as our final aligned model.
@@
Table 1: Basic information and number of sampled sentences for each language in the two language groups.
@@ 5 Results and Analysis
Starting with Mediterranean-Amharic-Farsi, we observe that the post-trained aligned model generally outperforms Glot500 on all downstream tasks. The SR-B task shows the most significant improvement, with the aligned model achieving, on average, more than 20% higher accuracy than Glot500 for all source languages. For other tasks, the aligned model also demonstrates a consistent, albeit smaller, improvement, with Glot500 occasionally outperforming the post-trained aligned model for specific source-target language pairs. However, we observe a more mixed performance for the South+East Asian Languages, especially for sequence labeling tasks, namely NER and POS, with the aligned models performing worse than Glot500 when transferring from more than half of the source languages. We hypothesize that this performance drop is primarily due to the transliteration process, which loses the semantic and contextual nuances and induces more token-level ambiguity for most of the languages (Amrhein and Sennrich, 2020; Liu et al., 2024a). This ambiguity makes token-level alignment more difficult. In contrast, the aligned model achieves consistent improvements in NER and POS for Mediterranean-Amharic-Farsi, where less token-level ambiguity is introduced as the languages are originally written in phonetic scripts, to which Latin also belongs.
@@
We observe that the performance can vary significantly for both groups of languages depending on the source language used for transfer. This phenomenon occurs due to the script and language similarity between specific source and target languages. In general, transferring from in-group high-resource languages performs better than transferring from English. Taking the SR-B task for example, for Mediterranean-Amharic-Farsi, the best performance is achieved when transferring from Farsi. For the South+East Asian Languages, the best performance is achieved when transferring from Korean. The text classification tasks generally show less variation in performance, though transferring from Hebrew achieves the worst performance for Mediterranean-Amharic-Farsi in Taxi1500. For the NER and POS tasks, transferring from Arabic and Hebrew achieves the best performance for Mediterranean-Amharic-Farsi while transferring from Chinese achieves the best performance for both tasks in South+East Asian Languages. Nevertheless, comparing our aligned models against Glot500, the performance generally improves for most source languages. This indicates that our proposed transliteration-based post-training method effectively improves the alignment between related languages and further boosts performance when a proper source language is chosen.
@@ 5.1 Ablation Study
The lone MLM objective already provides some slight improvement over the base Glot500 model. We hypothesize this is due to the benefit of specializing the model to a small group of languages. When the sequence-level alignment (SEQ) objective is included, the performance is generally further improved compared with the MLM variant, especially for the retrieval task. This is not surprising as the SEQ improves the sequence-level alignment. When instead the token-level alignment (TLM) objective is included, there is a slight improvement in the performance for most tasks compared to the MLM+SEQ objective, except for the retrieval task. We also observe that text classification tasks show the least variation in performance for most source languages. This is especially the case for SIB200, which seems to be the easiest task for the models.
Our whole objective (MLM+SEQ+TLM) generally performs better than the other models by combining sequence and token-level alignment benefits. For Mediterranean-Amharic-Farsi, though MLM+SEQ+TLM performs on par with the MLM+SEQ objective on NER, it achieves better performance on POS. Similarly, our complete training objective outperforms MLM+SEQ on both NER and POS for South+East Asian Languages. When comparing the performance of the full training objective against MLM+TLM, we observe an equal performance across different tasks, except for the retrieval task, where the full training objective noticeably outperforms MLM+TLM. Even though SEQ objective is the most critical for improving the retrieval task, we also observe a performance increase in the retrieval task when TLM is included. This indicates that our post-training alignment method, with TLM, effectively improves both the sequence and token-level alignment.

### v2

@@ 1 Introduction
Many post-training alignment strategies use objectives that rely on bilingual dictionaries or parallel data to align the representations of mPLMs (Cao et al., 2020; Wang et al., 2020; Schuster et al., 2019; Pan et al., 2021). However, dictionaries and parallel corpora are often limited in the data scale or the number of languages they cover (Artetxe et al., 2020b), which might be impractical for building strong supervision signals for many low-resource languages. Another alternative that improves cross-lingual alignment is to use transliteration (a process of converting the text of a language from one script to another). Transliteration can improve the lexical overlap, especially for related languages (Moosa et al., 2023). Unlike translation, transliterations can be obtained nearly for free using well-performing rule-based tools (Hermjakob et al., 2018). Therefore, several works have shown improvements in cross-lingual transfer by pre-training or fine-tuning models with data transliterated into a common script (Murikinati et al., 2020; Muller et al., 2021; Purkayastha et al., 2023; Moosa et al., 2023). However, these methods require the model to use a single common script, which is restrictive for many tasks as the transliteration process can be lossy and non-invertible.
@@
Recently, Liu et al. (2024a) proposed a sequence-level contrastive learning objective to improve the alignment across different scripts at a large scale (for more than 500 languages), using both original script sentences and their Latin transliterations, validated by English-centric cross-lingual zero-shot transfer evaluations. However, there are three major limitations in their setup. First, the contrastive objective only manipulates the sequence-level representations in the middle layer, which does not directly contribute to better alignment in the token-level space. Second, not every language pair has extensive lexical overlap that can boost cross-lingual transfer: transliteration-based alignment makes more sense for mutually influenced languages. Lastly, English alone as a transfer source language does not fully exploit the alignment benefit, as it does not have the most lexical overlap with other languages.
@@
To this end, we propose a new transliteration-based post-training alignment method that works on both sequence and token levels. Our method does not rely on parallel data. Instead, similar to Liu et al. (2024a), we use the monolingual data in their original scripts and their Latin transliteration obtained by using Uroman (Hermjakob et al., 2018). We investigate the impact of the strategy by focusing on two groups of languages: Mediterranean-Amharic-Farsi and South+East Asian Languages, described more in detail in Section 4.2. The languages in each group share areal features but differ in scripts. Some languages are closely related as members of the same language family (e.g., Semitic and Sino-Tibetan). Additionally, languages in each group have extensive lexical overlap due to historical contact and geographical proximity (e.g., Chinese and Korean or Turkish and Arabic). As these languages are written in different scripts, transliteration can help to better exploit the shared linguistic properties and thus improve the cross-lingual transfer.
@@
We leverage our method to post-train Glot500 (ImaniGooghari et al., 2023) (a continually pre-trained model from XLM-R on more than 500 languages) on the selected language groups and evaluate the zero-shot cross-lingual transfer performance on three types of downstream tasks: sentence retrieval, text classification, and sequence labeling. The evaluation is done with English as the source language and three other source languages of different scripts in each group. We show that our method consistently improves the downstream task performance across different languages and scripts. Moreover, our method further boosts the transfer performance when better source languages are chosen, as the performance depends on the degree of alignment between the source and target languages – precisely the alignment our approach boosts.
@@ 2 Related Work
Many recent works have proposed pre-training or fine-tuning alignment methods to improve cross-lingual transfer in mPLMs. Cao et al. (2020) proposed a fine-tuning embedding alignment objective between word pairs procured in an unsupervised fashion from parallel data using statistical word alignment models (Dyer et al., 2013). Chaudhary et al. (2020) improved alignment during pre-training by using bilingual dictionaries to replace words in original sentences with translations in other languages. Similarly, Tang et al. (2022) used bilingual dictionaries to explicitly align the embeddings of the same words in different languages during pre-training. Wei et al. (2021) proposed a hierarchical contrastive learning pre-training method, which uses parallel data to align representations at the word and sentence levels. Likewise, Hu et al. (2021) proposed a pre-training method with explicit alignment signals from parallel data that encourages symmetry at both word and sentence levels. Pan et al. (2021) combined contrastive learning with translation language modeling (Conneau and Lample, 2019) as a post-training alignment method that uses parallel data as well. While these methods have shown improvements in cross-lingual transfer, they have the limitation of requiring parallel data or bilingual dictionaries, which may be hard to acquire for many low-resource languages.
@@
Transliteration is a process of converting text from one language script to another (Wellisch et al., 1978). This process does not involve translating meanings but rather represents the original symbols as closely as possible in the target script. Different works have proposed transliteration-based methods to address the script barrier problem in multilingual models. Murikinati et al. (2020) used transliteration to a common script to improve cross-lingual morphological inflection. Khemchandani et al. (2021) exploited language relatedness between Indian languages and leveraged transliteration to a common script to adapt multilingual models to low-resource languages. Muller et al. (2021) analyzed the behavior of multilingual models on unseen languages and found that languages written in different scripts do not benefit from transfer learning. They proposed transliteration to the high-resource source language script to address the script barrier. Purkayastha et al. (2023) showed that fine-tuning multilingual models on data transliterated into Latin script improves cross-lingual transfer for low-resource languages. Similarly, Moosa et al. (2023) pre-trained models from scratch on data transliterated into a common script for the Indic languages and showed improvements in cross-lingual transfer. Our work is most related to TransliCo proposed by Liu et al. (2024a), where a sequence-level contrastive learning objective is used to encourage alignment across different scripts without restricting the models to a common script, using original script sentences and their Latin transliterations. However, their English-centric evaluation setup limits the ability to fully reveal the impact of transliteration. This paper systematically explores how transliteration-based alignment enhances transfer performance by using various source languages.
@@ 3 Methods
We present a transliteration-based post-training alignment method that can be used to fine-tune existing encoder-only mPLMs for improved alignment across languages using different scripts, boosting cross-lingual transfer performance. Our method consists of three objectives: masked language modeling, sentence-level alignment, and token-level alignment. All objectives are trained on combined original and transliterated data. The transliterated data is obtained by converting the original data into Latin script. The transliteration process uses Uroman (Hermjakob et al., 2018), a rule-based system that can convert nearly all character sets into a common Latin script. The overall method is illustrated in Figure 1, and we introduce the three objectives in detail below.
@@ 3.1 Masked Language Modeling
where $\mathcal{M}$ is the set of masked positions in the input sentence $X_{i}$ (either $X^{orig}_{i}$ or $X^{latn}_{i}$) and $p_{\text{MLM}}(X_{i,m}|\boldsymbol{h}_{i,m})$ is the probability of predicting token $X_{i,m}$ given $\boldsymbol{h}_{i,m}$, the final contextualized representation at the position $m$ in the $i$th sequence. The probability is computed by an MLM head. Fine-tuning with MLM on the original data is necessary to preserve the model’s knowledge. On the other hand, the mPLM has minimal knowledge about the transliterated data, which makes the MLM objective on transliterated data crucial for learning useful cross-script representations. We refer to the MLM objective for the original data (resp. transliterated data) as $\mathcal{L}_{\text{MLM}}^{orig}$ (resp. $\mathcal{L}_{\text{MLM}}^{latn}$).
@@ 3.2 Sentence-Level Alignment
where $\text{NEG}=\sum_{(X,X^{-})\in B^{-}}e^{\text{sim}(f(X),f(X^{-}))/\tau}$, $f$ is defined as mean pooling over the 8th layer output contextualized embeddings (ignoring the special tokens’ output except for [mask] token), sim is the dot product, and $\tau$ is the temperature set to $1$.
@@ 3.3 Token-Level Alignment
The sentence-level alignment objective helps the model to learn similar sentence representations for the original and transliterated sequences. This is useful for improving performance in sentence-level downstream tasks like sentence retrieval or classification. However, this objective manipulates the output of a middle layer, which does not directly contribute to better alignment in the token-level space. For token-level tasks like NER and POS tagging, alignment at the token level might be more beneficial. Therefore, we propose a token-level alignment objective that further encourages the model to align the representations of the original and transliterated words. We adapt the translation language modeling objective introduced by  Conneau and Lample (2019), which is equivalent to applying the MLM objective on a concatenated bilingual sentence pair. Specifically, given a sentence pair $(X^{orig}_{i},X^{latn}_{i})$, we apply the MLM objective on the concatenated sequence $X^{orig}_{i}\oplus X^{latn}_{i}$ or $X^{latn}_{i}\oplus X^{orig}_{i}$, where the concatenation order is randomly chosen during training. The intuition is that, to predict a token masked in the original sentence, the model can either attend to surrounding tokens in the original script or their transliterations and vice versa. This encourages the model to align the representations in the original and the Latin script. We refer to this objective as transliteration language modeling (TLM) and the loss as $\mathcal{L}_{\text{TLM}}$.
@@ 4.2 Languages, Data and Models
The two language groups are named as Mediterranean-Amharic-Farsi and South+East Asian Languages. We visualize each group’s geographical distribution of the selected languages in Figure 2. The languages within each group are spoken in adjacent areas, and there is a long history of linguistic influence between them. For example, Arabic has had extensive contact with languages such as Turkish and Persian (Versteegh, 2001). The data for each language is sampled from the Glot500-c training dataset (ImaniGooghari et al., 2023). We sample 10% of the available data for each language, or a minimum of 10k sentences, whichever is larger. The data is then transliterated into Latin script using Uroman (Hermjakob et al., 2018). Table 1 shows each group’s languages and the number of sampled sentences. In total, Mediterranean-Amharic-Farsi consists of 10 languages, 5 scripts, and around 16M sentences, while South+East Asian Languages consists of 10 languages, 7 scripts, and around 4M sentences. We fine-tune Glot500 using our alignment method on each group separately. We then select the best checkpoint for each group by validating the checkpoints’ performance on the Tatoeba (Artetxe and Schwenk, 2019) sentence retrieval dataset, which contains 1000 English-aligned sentences. We compute the top-10 retrieval accuracy based on the cosine similarity of the averaged 8th-layer contextual embeddings. The best checkpoint for each group is regarded as our final aligned model.
@@
Table 1: Basic information and number of sampled sentences for each language in the two groups.
@@ 5 Results and Analysis
Starting with Mediterranean-Amharic-Farsi, we observe that the post-trained aligned model generally outperforms Glot500 on all downstream tasks. The SR-B task shows the most significant improvement, with the aligned model achieving, on average, more than 20% higher accuracy than Glot500 for all source languages. For other tasks, the aligned model also demonstrates a consistent, albeit smaller, improvement, with Glot500 occasionally outperforming the post-trained aligned model for specific source-target language pairs. However, we observe a more mixed performance for the South+East Asian Languages, especially for the sequence labeling tasks, with the aligned models performing worse than Glot500 when transferring from more than half of the source languages. We hypothesize that this performance drop is primarily due to the transliteration process, which loses the semantic and contextual nuances and induces more token-level ambiguity for most of the languages (Amrhein and Sennrich, 2020; Liu et al., 2024a). This ambiguity makes token-level alignment more difficult. In contrast, the aligned model achieves consistent improvements in NER and POS for Mediterranean-Amharic-Farsi, where less token-level ambiguity is introduced as the languages are originally written in phonetic scripts, to which Latin also belongs.
@@
We observe that the performance can vary significantly for both groups depending on the source language used for transfer. This phenomenon occurs due to the script and language similarity between specific source and target languages. In general, transferring from in-group high-resource languages performs better than transferring from English. Taking the SR-B task, for example, for Mediterranean-Amharic-Farsi, the best performance is achieved when transferring from Farsi. For the South+East Asian Languages, the best performance is achieved when transferring from Korean. The text classification tasks generally show less variation in performance, though transferring from Hebrew achieves the worst performance for Mediterranean-Amharic-Farsi in Taxi1500. For the NER and POS tasks, transferring from Arabic and Hebrew achieves the best performance for Mediterranean-Amharic-Farsi, while transferring from Chinese achieves the best performance for both tasks in South+East Asian Languages. Nevertheless, comparing our aligned models against Glot500, the performance generally improves for most source languages. This indicates that our proposed transliteration-based post-training method effectively improves the alignment between related languages and further boosts performance when a proper source language is chosen.
@@ 5.1 Ablation Study
The lone MLM objective already slightly improves the base Glot500 model. We hypothesize this is due to the benefit of specializing the model to a small group of languages. When the sequence-level alignment (SEQ) objective is included, the performance is generally further improved compared with the MLM variant, especially for the retrieval task. This is not surprising as the SEQ improves the sequence-level alignment. When instead the token-level alignment (TLM) objective is included, there is a slight improvement in the performance for most tasks compared to the MLM+SEQ objective, except for the retrieval task. We also observe that text classification tasks show the least variation in performance for most source languages. This is especially the case for SIB200, which seems to be the most straightforward task for the models.
Our whole objective (MLM+SEQ+TLM) generally performs better than the other models by combining sequence and token-level alignment benefits. For Mediterranean-Amharic-Farsi, though MLM+SEQ+TLM performs on par with the MLM+SEQ objective on NER, it achieves better performance on POS. Similarly, our complete training objective outperforms MLM+SEQ on both NER and POS for South+East Asian Languages. When comparing the performance of the full training objective against MLM+TLM, we observe an equal performance across different tasks, except for the retrieval task, where the full training objective noticeably outperforms MLM+TLM. Even though the SEQ objective is the most critical for improving the retrieval task, we also observe a performance increase in the retrieval task when TLM is included. This indicates that our post-training alignment method, with TLM, effectively improves both the sequence and token-level alignment.